three times the difference of 18 subtracted from a certain number is 5 less than the number increased by 15. Find the number.
Found 2 solutions by Timnewman, lwsshak3:
Answer by Timnewman(323)

Hi dear,
Let the number be x
Three times the difference of 8 subtracted from a number is written as;3(x-8).
Five less than the number is increased by 15 is (x+15)-5.
But,
3(x-8)=x+15-5
Solving the resulting equation,
3x-24=x+10
3x-x=10+24
2x=34
x=17
The number is 17

Answer by lwsshak3(11628)

three times the difference of 18 subtracted from a certain number is 5 less than the number increased by 15. Find the number.
let x=number
3(x-18 )=(x+15)-5
3x-54+5=x+15
2x=44
x=22
number is 22